Suppose the demand equation for computers by Teetan Ltd for the year 2017 is given by Qd= 1200-P and the supply equation is given by Qs= 120+3P. Find equilibrium price and analyse what would be the excess demand or supply if price changes to Rs 400 and Rs 120
1
Expert's answer
2020-12-01T10:07:36-0500

At equilibrium Qd = Qs

1200-p = 120+3p

1200-120 = 3p+ p

1080 = 4p

P = 1080/4

P = 230

So equilibrium price = 230

If price changes to Rs 400 and Rs 120

Qd = Qs

Qd = 1200-p

Excess demand

1200- 400

= 800

Or

1200-120

= 1080